The Wildcat Oil Company is trying to decide whether to lease or buy a new computer-assisted drilling system for its oil exploration business. Management has decided that it must use the system to stay competitive; it will provide $2.4 million in annual pretax cost savings. The system costs $7.1 million and will be depreciated straight-line to zero over five years. Wildcat's tax rate is 23 percent, and the firm can borrow at 8 percent. Lambert Leasing Company has offered to lease the drilling equipment to Wildcat for payments of $1.63 million per year. Lambert's policy is to require its lessees to make payments at the start of the year. Suppose Lambert requires Wildcat to pay a $475,000 security deposit at the inception of the lease. Calculate the NAL with the security deposit. (A negative answer should be indicated by a minus sign. Do not round intermediate calculations and enter your answer in dollars, not millions, rounded to 2 decimal places, e.g., 1,234,567.89.)
Answer #1

A C D E F G H 1 B Buy Year Cost Annual post tax savings Dep tax savings 0 1 2 3 4 5 -7100000 1848000 1848000 1848000 1848000

D E F G H I 1 2 3 4 5 Ham to do BC Buy Year o Cost =-7.1*10^6 Annual post tax savings Dep tax savings = 2.4*10^6 *(1-23%) =2.